16) Calculate the working capital amount given the following data: Be sure to properly label your answer. Show all work.

Cash $10,000

A/R 20,000

           Inventory 10,000

PPE (LTA) 50,000

Notes Payable (CL) 10,000

Mortgage Payable (LTL) 40,000

Retained Earnings 40,000

Rent expense 1,000

Car expense 500

Shipping expense 100

* LTL = long term liability, CL = current liability, LTA= Long term asset



17) Calculate the current ratio using the data in problem 16. Be sure to properly label your answer. Show all your work.
Answer #1

16.

Working capital is the difference between current assets and current liabilities. Working capital is essential running capital to perform daily business operations.

Current assets = Cash + A/R + Receivables = 10,000 + 20,000 + 10,000 = $ 40,000

Current liabilities = Notes payable = $ 10,000

Working capital = Current assets – current liabilities = 40,000 – 10,000 = $ 30,000

17.

Current ratio is liquidity ratio depicting the ability of company to meet its short-term obligations within 1 year. To compute, current ratio, current assets is divided by current liabilities.

Using data from 16,

Current assets = $ 40,000

Current liabilities = $ 10,000

Current ratio = current assets/current liabilities = 40,000/10,000 = 4

kindly upvote